the manager of farmers delight bought a load of watermelons for $750 and priced the watermelons so that he would make $2 on each melon. when all but 100 of the melons had been sold, he broke even. how many did he originally buy?
Answer by 314(4)

x=number of watermelons
2(x-100)=750 we divide with 2
x-100=375
x=375+100
x=475 watermelons